OpenStudy (anonymous):

consider the line y= 10-2x amd the function y(x)=2x^3 - 3x^2 -14x +3. determine the coordinates of the point(s) where the line and the graph of the function are tangent. THANK YOU!!!

OpenStudy (tkhunny):

The slope of the line is what? The general expression for the slope of the tangent is what?

OpenStudy (anonymous):

slope of line =2 and i didnt get the second part

OpenStudy (tkhunny):

Have you considered the 1st Derivative, y'(x)?

OpenStudy (anonymous):

oh, that would be 6x^2 -6x-14, right?

OpenStudy (tkhunny):

Yes. Where does y'(x) = -2? BTW, the slope of the line is -2, not just 2.

OpenStudy (anonymous):

i'l try to work out the final answer and i'll post it

OpenStudy (anonymous):

i found\[0.5+(\sqrt{420}/12)\] and \[0.5-(\sqrt{420}/12)\] not sure how to simplify the sqrt of 420/12

OpenStudy (tkhunny):

Oh, how I wish you had shown your work. Something went seriously wrong. You should be trying to solve: \(6x^{2} - 6x - 14 = -2\) Is that where you started?

OpenStudy (anonymous):

yeah sry, i sent -2 to the other side of the equation instead of 2. so, D= 324 \[0.5-( \sqrt{324}/12), 0.5+( \sqrt{324}/12)\]

OpenStudy (tkhunny):

\(\sqrt{324} = 18\)

OpenStudy (anonymous):

x=-1 x=2 am i supposed to use the -1?

OpenStudy (tkhunny):

Why would you ignore it? There are two solutions.